OpenStudy (anonymous):

@acxbox22 Can you go over my work to see if its right? I'm not positive It's correct.

OpenStudy (anonymous):

can you show it?

OpenStudy (anonymous):

This was my assignment for question 2 I got f(3)=5. Is that correct?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Alright thank you :)

OpenStudy (anonymous):

For question 3 would this be correct @zimmah You would substitute the f(x) with Y. Your equation would then look like y=4x-7. After that you would switch out the Y and X. Then the function would look like x=4y-7 . Then add the 7 on both sides. The seven then cancels out. You should get x-4=11y, Then you should divide the 11 by both sides. Your answer should be (fx)=1/11x-4/11.

OpenStudy (acxbox22):

perfect...looks good

OpenStudy (acxbox22):

except x=4y-7 x+7=4y x/4 +7/4=y

OpenStudy (acxbox22):

the 7 doesnt cancel out

OpenStudy (anonymous):

So you wouldn't cancel out the seven at all?

OpenStudy (acxbox22):

no you would just add the 7 on both sides to isolate 4y

OpenStudy (anonymous):

Can you help me with my last problem. not give me the answer but explain it? given the equation g(x)=-2x+6 how can we evaluate f(g(x))?

OpenStudy (acxbox22):

you know how gx = -2x+6 substitute that for x in function f(x)

OpenStudy (acxbox22):

so f(g(x))=4(-2x+6)+6

OpenStudy (anonymous):

Okay okay. I got it. Thank you

OpenStudy (acxbox22):

no problem